For an APR of 10% per year, if the interest is compounded weekly, determine the nominal rate per (a) 6 months and (b) 2 years.

The nominal interest rate per 6 months is %.

The nominal interest rate per 2 years is %.
Answer #1

APR per year = 10%

As, rate is compounded weekly, therefore, monthly interest rate = 10/52 = 0.19%

As, Nominal interest rate = ((1+r)n - 1), where r is interest rate and n is number of periods

a) Now, Nominal interest rate for 6 months or 26 weeks = (1+0.0019)26 - 1, as there are 26 weeks

Therefore, Nominal interest rate for 6 months = 5.06%

b) And, Nominal interest rate for 2 years or 104 weeks = (1+0.0019)104 - 1, as there are 104 weeks in 2 years

Therefore, Nominal interest rate for 2 years = 21.82%